OpenStudy (anonymous):

you and some friends are planning a snowboarding holiday, to take place on December 28. the snow depths at three nearby ski resorts on this date in four recent years are shown in the table below, which you have complied by gathering data from the internet Dec.28, 2000: Gravity Gulch 74, Mt. Thrasher 81, Xtreme Terror 54. Dec.28, 2001: Gravity Gulch 37, Mt. Thrasher 34, Xtreme 47. Dec.28,2002: Gravity Gulch 56, Mt. Thrasher 56, Xtreme Terror 80. Dec.28, 200: Gravity Gulch 72, Mt. Thrasher 73, Xtreme Terror.

OpenStudy (anonymous):

2003*

OpenStudy (anonymous):

For example, on December 28, 2002, Mt. Thrasher had 56 inches of snow. Analyze the data and predict which resort will have the best snow for your holiday

OpenStudy (anonymous):

Just take the average of each one (ex:Gravity Gulch) and then see which one has the highest average?

OpenStudy (anonymous):

A.Mt. Thrasher B.Gravity Gulch C.Xtreme Terror

OpenStudy (anonymous):

For Gravity Gulch, average would be = (74+37+56+72)/4 = 59.75 For Mt. Thrasher, average = (81+34+56+73)/4 = 61 For Xtreme terror, the info for 2003 is missing. so it's the same method just take the average and choose the 1 with the higest average.